OpenStudy (anonymous):

How many and of what type are the solutions to x2 + 10x + 25 = 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

theres (x+5)(x+5) and there identical rational solutions

OpenStudy (anonymous):

so its two identical rational solutions.

OpenStudy (anonymous):

i was right jkkkk

OpenStudy (anonymous):

wait so you werent right?

OpenStudy (anonymous):

@MathDoodler what are you doing?

OpenStudy (anonymous):

To solve, use the following formula to find the discriminant \[ \sqrt{b^2-4ac} \] Plugin your values from x2 + 10x + 25 = 0 ax + bx + c = 0 \[ \sqrt{10^2-4(1)(25)} \] \[ \sqrt{100-100} \] \[ \sqrt{0} \] The discriminant = 0 Rule for discriminant If discriminant = 0 the equation has two equal solutions, a double root If discriminant > 0 the equation has two unequal real number solution If discriminant < 0 the equation has two complex number solutions so your answer is Two identical rational solutions

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

Playing COD and trying to type so that is what was taking me so long lol